A corporation creates a sinking fund in order to have $670,000 to replace some machinery in 8 years. How much should be placed in this account at the end of each month if the annual interest rate is 6.5% compounded monthly? (Round your answers to the nearest cent.)
$

How much interest would they earn over the life of the account?
$

Determine the value of the fund after 2, 4, and 6 years.

2 years $
4 years $
6 years $


How much interest was earned during the second month of the 7th year?
$ 5. + -16 points TEAFM2 F.3.026. My Notes A corporation creates a sinking fund in order to have $670,000 to replace some machi
Answer #1

1.
=670000*(6.5%/12)/((1+6.5%/12)^(12*8)-1)=5339.60918056332

2.
=670000-670000*(6.5%/12)/((1+6.5%/12)^(12*8)-1)*12*8=157397.518665921

3.
=670000*(6.5%/12)/((1+6.5%/12)^(12*8)-1)*1/(6.5%/12)*((1+6.5%/12)^(12*2)-1)=136459.643320511

4.
=670000*(6.5%/12)/((1+6.5%/12)^(12*8)-1)*1/(6.5%/12)*((1+6.5%/12)^(12*4)-1)=291809.249458764

5.
=670000*(6.5%/12)/((1+6.5%/12)^(12*8)-1)*1/(6.5%/12)*((1+6.5%/12)^(12*6)-1)=468663.735811384

6.
=670000*(6.5%/12)/((1+6.5%/12)^(12*8)-1)*1/(6.5%/12)*((1+6.5%/12)^(73)-1)*(6.5%/12)=2581.26884